PepsiPFU Commissioning Observing - 20131128 UT

Observers: JStorm, IIlyin, MWeber (LBTO), DMiller, JHill (Tucson)
Telescope Operator: DHuerta (LBTO)
AO Support:: DMiller (Tucson)
Telescope Support:: JLittle (LBTO)
Instrument Support:: JMorris (LBTO)
Engineering Support: DAshby (Tucson)

Summary

Closed for clouds at the beginning.

Then field aberration data to adjust rigid secondary at LFBG

PepsiPFU Commissioning for 3 hours starting at 03:20 UT. We found stars on both sides and were able to do basic guiding and offsetting with some wrinkles still to be worked out.

Details

Field Aberration Measurements at LFBG for Rigid Secondary

Initial Pointing Check in Binocular Mode

Default Pointing from 20131025 SX pointing model: SX IE=-49 CA=-40

Adjusted Both Sides to center DX star, but SX star not found at: SX IE=-33 CA=-26

Found the SX star at: SX IE=-30 CA=-64, but why was the pointing worse than last night?

Preset ACQUIRE to M5 star and when a star was found (IE = -49 to -30 and CA = -40 to -64). Coma was noticeable but not to large. We decide to not make a move and collect field data

Collimation on BS9101

Field Data on Stone_P1

Collect LFBG field on Stone_P1. Data in dir obs2:/home/dlmiller/data/20131128/field

tuco:/work/field/lfbg/20131128

 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism
  
                  lfbg
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  17.5", Delta IE = -17.5"
         Tilt = -20.8", Delta CA =  20.8"
  
 
  => Field Data Results
             Measured bi-nodal astigmatism
  
                  lfbg
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  25.9", Delta IE = -25.9"
         Tilt = -36.7", Delta CA =  36.7"
  
  
                  lfbg
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  27.2", Delta IE = -27.2"
         Tilt = -28.7", Delta CA =  28.7"
 

02:55 Apply the above binodal corrections to SX M2: Tip=+25 Tilt=-30 dIE=-25 dCA=+30

New SX IE=-55 CA=-34

SX M2 Global: X=-2.926 Y=-0.872 Z=-3.900 RX=-519.6 RY=-423.6 (from 25 October)
SX M2 Manual: X=+1.532 Y=-1.277 Z=0.000 RX=+245.5 RY=+294.6 (from tonight's measured TT correction)
SX M2 Active: X=-0.848 Y=+2.029 Z=+0.212 RX=-211.4 RY=-84.3 (from active optics)

SX New Global: X=-2.242 Y=-0.120 Z=-3.688 RX=-485.5 RY=-213.3 (sum of the above)

These were installed in all the SX Gregorian Collimation models after the PEPSI work.

PepsiPFU commissioning on both sides

03:20 Authorize at PepsiPFU both left and right

03:30 both artificial stars were still in. Ilya removed them and we see a glint on SX but nothing on DX

03:33 David move telescope IE CA (both mode) so star is on acquisition image
SX IE=-35 CA=-36 including the tip-tilt for binodal correction.

03:43 John change IE by +5. Star move down and left (235 degrees). This is + Elevation

03:45 John change CA by +5. Star move down and right (315 degrees). This is + Azimuth

03:47 John defocus to adjust the Coma. Bright to upper right so in -Elevation direction. Send -3300 Z7 (~+2 mm in Y)

03:52 Az direction send +1000 Z8

03:54 iif_offset, 5, 0 DETXY. Star move down and to the right. +Az direction

03:55 SX Primary panic. Move to Zenith and set the SX Secondary. Jesper and Michael go to telescope and shine a light into PFU. We see the two pupils on lower left. Thus, guider is sensitive and PFU box shutter is open. Is Tertiary is incorrect position?

04:05 Set RIGHTIAA to 47.0. Restart PCS, put in IE, CA. Preset acquire.

04:06 iif_offset, 5, 0 DETXY. Star move straight down. Wrong sign

04:07 Set RIGHTIAA to -47.0. Set LEFTIAA to 47.0. Restart PCS, put in IE, CA. Preset acquire.
  • iif_offset, 5, 0, DETXY. Star moves to the right. Correct
  • iif_offset, 0, 5, DETXY. Star moves down. Incorrect.

04:10 Change LEFTXYFLIP from "false false" to "false true"
  • Restart PCS, put in IE, CA. Preset acquire.

04:28 Center star IE = -35, CA = -31

04:33 Binoc Preset BS9107 ACQUIRE. See star off top of LEFT, on left side of RIGHT

04:43 Test offsets in ROTATOR=POSITION mode (previous tests were in PARALLACTIC mode)
  • iif_offset, 5, 0, DETXY. Star moves up. Incorrect
  • iif_offset, 0, 5, DETXY. Star moves left. Incorrect

04:45 Set RIGHTIAA to -47.0. Set LEFTIAA to 47.0. Restart PCS, put in IE, CA. Preset acquire.

04:49 RIGHT offset
  • iif_offset, 5, 0, DETXY. Star moves up. Incorrect
  • iif_offset, 0, 5, DETXY. Star moves right. Incorrect

04:52 Change RIGHTXYFLIP from "false false" to "false true"

04:54 David Stop and Start PCS

04:58 John set LEFT IE, CA

04:59 preset LEFT GUIDE.
  • Star close on acquisition

05:05 Change GCS flips
  • flip_x from -1 to +1
  • flip_y from +1 to -1

05:11 Change GCS flips
  • flip_y from +1 to +1
  • star moves down and right

05:16 Change GCS flips
  • flip_y from +1 to +1
  • star moves down and right

05:22 RIGHT guiding is off by about +20 degrees (more CW direction)

05:28 gcs_set_hotspot, 340, 360, /RIGHT (from 340, 370)

05:30 gcs_set_hotspot, 360, 386, /RIGHT (from 370, 386)

05:36 preset LEFT ACTIVE
  • WFSC image too dim error from GCS. Did not process wfsc image
  • images 80, 81

05:35

05:39 binoc preset to BS9107, TS=1 (10 mag)
  • clouds

RIGHT WFSC Rotation Angle Data

image 09 start

z4? +1000, image 12 13
z4? -1000, 15
z0 0 16

z5 +1000, image 17, 18
z5 0
z5 +1000 image 20
z5 0 21

z6 +1000 image 22, 23

z9 +500 image 28
z9 0

z10 +500 image 30, 31

LEFT WFSC Rotation Angle Data

138 start

z5 +1000 139, 140, 141
z5 0

z6 +1000 143

z9 +500 147, 148

z10 +500 149

-- DougMiller - 27 Nov 2013
Topic revision: r5 - 28 Nov 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback